OpenStudy (anonymous):

Help with Ellipses? :c Find the center, vertices, and foci of the ellipse with equation 2x^2 + 6y^2 = 12

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@phi

OpenStudy (phi):

put the equation into standard form

OpenStudy (phi):

standard form looks like this \[ \frac{(x-h)^2 }{a^2} + \frac{(y-k)^2}{b^2 }=1 \] the very first step is make the right side = 1 (by dividing both sides by 12) can you do that ?